# Ambipolar doping of a charge-transfer insulator in the Emery model

### v1

@@ X Conclusions
In this work, we have considered similarities and differences of doping with electrons and holes a charge-transfer insulator realised by the Emery model in the normal state, solved with CDMFT. We have addressed two main questions: (i) what is the nature of the doping driven metal to charge-transfer insulator transition and (ii) how the orbital character of the dopant carriers evolves with either doping paths. This work expands Ref. [44], which has used the same methodology of this work and has focused on the hole-doping case only.
Regarding the first question about the nature of the metal to charge-transfer insulator transition, we find a purely electronic two-stage doping driven transition upon either electron or hole doping, i.e. irrespective of the orbital character of the dopants. Above a critical threshold of $U_{d}$ for opening a charge-transfer gap at $\delta=0$, there is a continuous transition between a charge-transfer insulator and a strongly correlated pseudogap at $\delta=0$, followed by a discontinuous transition between a strongly correlated pseudogap and a correlated metal at finite doping. The pseudogap to metal transition ends in a critical point at finite temperature, from which emerges a supercritical crossover marking an anomalous enhancement in the charge compressibility. This two-stage electronic transition occurs upon either electron or hole doping (see $T-\delta$ phase diagram in Fig. 5). By varying the strength of the electron-electron interaction $U_{d}$ on the copper orbital, we find that the pseudogap to metal transition on either doping path is connected to the metal to charge-transfer insulator transition at zero doping (see $U_{d}-\delta$ phase diagram in Fig. 11). Thus, a surface of first order transition in the $U_{d}-T-\delta$ space crosses zero doping and progressively extends to larger dopings with increasing $U_{d}$.
From these findings we identify three main conclusions. (i) There is no qualitative electron-hole asymmetry in the nature of the doping driven transition of a charge-transfer insulator in the normal state. A quantitative electron-doping asymmetry is however found on the doping level at which this transition occurs and on the temperature at which this transition ends. (ii) In agreement with the results of Ref. [44], the pseudogap is a strongly correlated phase that only emerges for $U_{d}$ larger than the critical threshold to open a charge-transfer insulator at $\delta=0$ (one hole per CuO2 unit cell), hence the name ‘strongly correlated pseudogap’. (iii) Beyond the results of Ref. [44], the pseudogap on either doping path has a common origin in Mott physics in two dimensions, that is, it is a robust phenomenon emerging from a doped correlated insulator, regardless of the (hole vs electron) nature of the dopants.
@@
Taken together, our results show that ambipolar doping of a charge-transfer insulator produces an asymmetry in the orbital character of the dopants, but it is not sufficient to produce an asymmetry in the qualitative features of the normal state doping driven transition. Hence, our findings rule out that the origin of the observed normal state electron-hole asymmetry in cuprates lies solely in doping with electrons and holes a charge-transfer insulator. In the Discussion Sec. IX, we commented of the implications for actual hole- and electron-doped cuprates.

### v2

@@ X Conclusions
In this work, we have considered similarities and differences of doping with electrons and holes a charge-transfer insulator realised by the Emery model in the normal state, solved with CDMFT. We have addressed two main questions: (i) what is the nature of the doping-driven metal to charge-transfer insulator transition and (ii) how the orbital character of the dopant carriers evolves with either doping paths. This work expands Ref. [44], which has used the same methodology of this work and has focused on the hole-doping case only.
Regarding the first question about the nature of the metal to charge-transfer insulator transition, we find a purely electronic two-stage doping-driven transition upon either electron or hole doping, i.e. irrespective of the orbital character of the dopants. Above a critical threshold of $U_{d}$ for opening a charge-transfer gap at $\delta=0$, there is a continuous transition between a charge-transfer insulator and a strongly correlated pseudogap at $\delta=0$, followed by a discontinuous transition between a strongly correlated pseudogap and a correlated metal at finite doping. The pseudogap to metal transition ends in a critical point at finite temperature, from which emerges a supercritical crossover marking an anomalous enhancement in the charge compressibility. This two-stage electronic transition occurs upon either electron or hole doping (see $T-\delta$ phase diagram in Fig. 5). By varying the strength of the electron-electron interaction $U_{d}$ on the copper orbital, we find that the pseudogap to metal transition on either doping path is connected to the metal to charge-transfer insulator transition at zero doping (see $U_{d}-\delta$ phase diagram in Fig. 12). Thus, a surface of first order transition in the $U_{d}-T-\delta$ space crosses zero doping and progressively extends to larger dopings with increasing $U_{d}$.
From these findings we identify three main conclusions. (i) There is no qualitative electron-hole asymmetry in the nature of the doping-driven transition of a charge-transfer insulator in the normal state. A quantitative electron-doping asymmetry is however found on the doping level at which this transition occurs and on the temperature at which this transition ends. (ii) In agreement with the results of Ref. [44], the pseudogap is a strongly correlated phase that only emerges for $U_{d}$ larger than the critical threshold to open a charge-transfer insulator at $\delta=0$ (one hole per CuO2 unit cell), hence the name ‘strongly correlated pseudogap’. (iii) Beyond the results of Ref. [44], the pseudogap on either doping path has a common origin in Mott physics in two dimensions, that is, it is a robust phenomenon emerging from a doped correlated insulator, regardless of the (hole vs electron) nature of the dopants.
@@
Taken together, our results show that ambipolar doping of a charge-transfer insulator produces an asymmetry in the orbital character of the dopants, but it is not sufficient to produce an asymmetry in the qualitative features of the normal state doping-driven transition. Hence, our findings rule out that the origin of the observed normal state electron-hole asymmetry in cuprates lies solely in doping with electrons and holes a charge-transfer insulator. In the Discussion Sec. IX, we commented of the implications for actual hole- and electron-doped cuprates.